# Image Embedding Sampling Method for Diverse Captioning

### v1

@@ Abstract
Image Captioning for state-of-the-art VLMs has significantly improved over time; however, this comes at the cost of increased computational complexity, making them less accessible for resource-constrained applications such as mobile devices and assistive technologies. Alternatively, smaller VLMs prioritize high-level scene descriptions, overlooking finer details that contribute to a richer understanding of an image. In this paper, we introduce a training-free framework that enhances caption diversity and informativeness by explicitly attending to distinct image regions using a comparably small VLM, BLIP, as the backbone. Our approach leverages structured segmentation to produce hierarchical representations that capture both global and localized semantics. Without requiring additional model training, we demonstrate that our method allows smaller VLMs to achieve performance comparable to larger models in terms of image-caption alignment, semantic integrity, and diversity. We evaluate our framework on MSCOCO, Flickr30k, and Nocaps test datasets, achieving a Div-2 score of 0.735, 0.750, and 0.748 for each dataset respectively, while maintaining strong image-caption relevancy and semantic integrity with the human-annotated captions.
[…]
latex‘!h’ float specifier changed to ‘!ht’
@@
Sania Waheedthanks: Equal contribution   Na Min An11footnotemark: 1 University of Southhampton, KAIST sw1m24@soton.ac.uk, naminan@kaist.ac.kr
@@ 1 Introduction
Visual-Language Models (VLMs) have seen rapid advancements in image captioning, benefiting from increasingly sophisticated architectures and larger training datasets Alayrac et al. (2022); Li et al. (2022b); Radford et al. (2021a); Wang et al. (2022a). State-of-the-art large-scale models generate highly detailed and diverse captions, yet their extensive computational requirements can be prohibitive in resource-constrained settings. Conversely, smaller VLMs, while more efficient, often prioritize dominant visual elements and overlook fine-grained details, resulting in captions that lack the depth and specificity seen in human-generated captions Aneja et al. (2019a); Bianco et al. (2023); Chen et al. (2023); Yuksekgonul et al. (2022).
Inspired by previous work Ji et al. (2021); Shao et al. (2023); Shukor et al. (2022) that demonstrates the advantages of hierarchical approaches in image understanding, our method leverages structured segmentation to capture both global and regional aspects of an image. We sample segmentation-driven embeddings to explicitly attend to distinct image regions while preserving contextual relationships, generating captions at multiple levels of granularity. This approach enables smaller VLMs to achieve performance comparable to larger models in terms of caption diversity and image-caption alignment.
We validate our approach, namely, HBoP - Hierarchical Bags of Phrases, by evaluating generated captions for MSCOCO, Flickr30k, and Nocaps datasets on conventional diversity metrics such as mBLEU-4, n-gram diversity Aneja et al. (2019b), and newly presented pairwise cosine distance (PCD). Our findings show that structured caption generation effectively improves diversity while maintaining relevancy with images and human-generated captions (compare BLIP Li et al. (2022a), HBoP, and gold captions in Figure 1).
@@ 2 Related Works
Vision-language models have demonstrated remarkable capabilities in various multi-modal tasks, with caption generation being a primary benchmark for evaluating their performance. Models such as CLIP Radford et al. (2021a), Flamingo Alayrac et al. (2022), and BLIP-2 Li et al. (2023) leverage contrastive learning and large-scale pre-training to improve vision-language alignment. Although these approaches improve caption fluency and coherence, they often generate high-level scene descriptions without capturing fine-grained details, limiting their utility for applications requiring detailed image understanding. Traditional captioning models treat images holistically, often overlooking fine-grained hierarchical details Xu et al. (2021), unless trained with an explicit training objective of improving diversity, such as ModeCap Chen et al. (2022) which explores various modes in the training corpus and Seq-CVAE Aneja et al. (2019b) that uses latent variables to generate every word within a sentence.
Inspired by hierarchical representation techniques Ji et al. (2021); Shao et al. (2023); Shukor et al. (2022), our approach samples the latent image embeddings corresponding to structured segmentation to generate captions that reflect multiple levels of contextual details. PnP-VQA Tiong et al. (2022) proposes a similar methodology but focuses primarily on visual question answering. While PnP-VQA leverages image segmentation to generate captions, its sampling approach results in less meaningful captions due to its reliance on high-activation regions extracted using Grad-CAMs Selvaraju et al. (2017).
Figure 2: The proposed HBoP framework consists of three components: (1) Image Segmentation Module (ISM), (2) Hierarchical Composition Module (HCM), and (3) Image Captioning Module (ICM). HBoP controls caption granularity by selecting meaningful patch embeddings of varying sizes from the segmentation model.
@@ 3.1 Image Segmentation Module (ISM)
The first component of HBoP, ISM, selects patch embeddings ($E_{X}$) corresponding to image regions ($X=(X_{1},X_{2},...,X_{n})$) from the original image embeddings extracted using a Vision Transformer (ViT) Dosovitskiy et al. (2020) encoder. These regions are selected based on the segmentation masks produced by the state-of-the-art Segment Anything model (SAM) Kirillov et al. (2023). If we select a set of $p$ segmentation masks for the image, the resulting masks for the selected image regions would be: $M_{X}=\{\,M_{X_{1}},M_{X_{2}},...,M_{X_{p}}\,\}=\text{SAM}(X),X\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and channels of $X$.
@@ 3.2 Hierarchical Composition Module (HCM)
The global segmentation masks ($M_{G}$) are selected by choosing the top-$k$ (5 in our case) largest segmentation masks from $M_{X}$ after applying non-maximum suppression (NMS) Hosang et al. (2017):
$M_{G}$ =$=\{M_{g_{1}},M_{g_{2}},...,M_{n_{g}}\},$
[…]
@@
$M_{R}$ =$=\{M_{r_{1}},M_{r_{2}},...,M_{K}\},$
[…]
The hierarchical segmentation masks ($M_{g}$, $M_{r}$ and $M_{f}$) are used to extract relevant patch embeddings, $E_{g}$, $E_{r}$ and $E_{f}$ using $E_{X}$ from the first stage. We extract ($\odot$) the corresponding embeddings by concatenating the extracted patch embeddings of different levels. Thus, the final selected image embeddings can be categorized as:
$E_{G}$ =$=\{\,E_{g_{1}},E_{g_{2}},...,E_{g_{n_{g}}}\,\},E_{g_{i}}=E_{X}\odot M_{g_{i}}$
$E_{R}$ =$=\{\,E_{r_{1}},E_{r_{2}},...,E_{K}\,\},E_{r_{i}}=E_{X}\odot M_{r_{i}}$
$E_{F}$ =$=\{\,E_{f_{1}},E_{f_{2}},...,E_{n_{f}}\,\},E_{f_{i}}=E_{X}\odot M_{f_{i}}$
3.3 Image Captioning Module (ICM)
$\text{HBoP}_{G}$ =$=\{\,s_{g_{1}},...,s_{n_{g}}\,\},s_{g_{i}}=\text{BLIP}(E_{g_{i}})$
$\text{HBoP}_{R}$ =$=\{\,s_{r_{1}},...,s_{K}\,\},s_{r_{i}}=\text{BLIP}(E_{r_{i}})$
$\text{HBoP}_{F}$ =$=\{\,s_{f_{1}},...,s_{n_{f}}\,\},s_{f_{i}}=\text{BLIP}(E_{f_{i}})$
4 Results
@@
Encoder Param SBERT $\uparrow$ CLIP-S $\uparrow$ PCD mBLEU-4 $\downarrow$ Div-2 $\uparrow$ SBERT $\uparrow$ CLIP-S $\uparrow$ PCD mBLEU-4 $\downarrow$ Div-2 $\uparrow$
@@
Table 1: Relevancy and diversity scores across different model captions. HBoP achieves higher diversity scores with higher Div-2 and PCD values, and a lower mBLEU-4 score, on both the MSCOCO and Flickr30K datasets compared to smaller VLMs or models trained to enhance diversity. It also maintains comparable relevance scores (SBERT and CLIP-S). Simultaneously, HBoP scores competitive results compared to much larger VLMs with LLM encoders. The cells are colored based on comparison to HBoP, with blue indicating lower values and red indicating higher values.

### v2

@@ Abstract
Image Captioning for state-of-the-art VLMs has significantly improved over time; however, this comes at the cost of increased computational complexity, making them less accessible for resource-constrained applications such as mobile devices and assistive technologies. Alternatively, comparably smaller VLMs prioritize high-level scene descriptions, overlooking finer details that contribute to a richer understanding of an image. In this paper, we introduce a training-free framework that enhances caption diversity and informativeness by explicitly attending to distinct image regions using a comparably small VLM, BLIP, as the backbone. Our approach leverages structured segmentation to produce hierarchical representations that capture both global and localized semantics. Without requiring additional model training, we demonstrate that our method allows smaller VLMs to achieve performance comparable to larger models in terms of image-caption alignment, semantic integrity, and diversity. We evaluate our framework on MSCOCO, Flickr30k, and Nocaps test datasets, achieving a Div-2 score of 0.735, 0.750, and 0.748 for each dataset, respectively, while maintaining strong image-caption relevancy and semantic integrity with the human-annotated captions. Our code is available at https://github.com/xfactlab/HBoP.
@@
Sania Waheed [^1] Equal contribution. [^2] Work done while at KAIST. University of Southhampton sw1m24@soton.ac.uk          Na Min An [^1] Equal contribution. KAIST AI naminan@kaist.ac.kr
@@ 1 Introduction
Visual-Language Models (VLMs) have seen rapid advancements in image captioning, benefiting from increasingly sophisticated architectures and larger training datasets Alayrac et al. (2022); Li et al. (2022b); Radford et al. (2021a); Wang et al. (2022a). State-of-the-art large-scale models generate highly detailed and diverse captions, yet their extensive computational requirements can be prohibitive in resource-constrained settings. Conversely, smaller VLMs, while more efficient, often prioritize dominant visual elements and overlook fine-grained details, resulting in captions that lack the depth and specificity seen in human-generated captions Aneja et al. (2019a); Bianco et al. (2023); Chen et al. (2023); Yuksekgonul et al. (2022); An et al. (2025).
Inspired by previous work Ji et al. (2021); Shao et al. (2023); Shukor et al. (2022) that demonstrates the advantages of hierarchical approaches in image understanding, our method leverages structured segmentation to capture both global and regional aspects of an image. We sample segmentation-driven embeddings from the last layer of the visual encoder, where self-attention mechanisms have already propagated information across the entire image. This allows the model to explicitly attend to distinct local image regions while preserving contextual relationships, generating captions at multiple levels of granularity. This approach offers an efficient alternative to enhancing caption diversity in smaller VLMs without LLMs, achieving performance comparable to larger LLM-based models in terms of caption diversity and image-caption alignment.
We validate our approach, namely, HBoP - Hierarchical Bags of Phrases, by evaluating generated captions for MSCOCO Lin et al. (2014), Flickr30k Young et al. (2014), and Nocaps Agrawal et al. (2019) datasets on conventional diversity metrics such as mBLEU-4, n-gram diversity Aneja et al. (2019b), and newly presented pairwise cosine distance (PCD). Our findings show that structured caption generation effectively improves diversity while maintaining relevancy with images and human-generated captions (compare BLIP Li et al. (2022a), HBoP, and gold captions in Figure 1).
@@ 2 Related Works
Vision-language models have shown strong performance in multimodal tasks, with caption generation as a key benchmark. Models like CLIP Radford et al. (2021a), Flamingo Alayrac et al. (2022), and BLIP-2 Li et al. (2023) use contrastive learning and large-scale pre-training to enhance vision-language alignment. However, they often produce high-level scene descriptions, missing fine-grained details needed for detailed image understanding. Traditional captioning approaches treat images holistically, overlooking hierarchical details Xu et al. (2021), unless explicitly trained for diversity, as in ModeCap Chen et al. (2022) and Seq-CVAE Aneja et al. (2019b).
Inspired by hierarchical representation techniques Ji et al. (2021); Shao et al. (2023); Shukor et al. (2022), our approach samples latent image embeddings from structured segmentation to generate multi-level captions. This aligns with recent region-based methods using SAM Shlapentokh-Rothman et al. (2024) and studies on caption quality focused on informational sufficiency, minimal redundancy, and human comprehensibility Chen et al. (2024). Our evaluation metrics reflect these aspects: CLIP score for informational sufficiency, mBLEU and Div-2 for redundancy, and SBERT for comprehensibility.
Figure 2: The proposed HBoP framework consists of three components: (1) Image Segmentation Module (ISM), (2) Hierarchical Composition Module (HCM), and (3) Image Captioning Module (ICM). HBoP controls caption granularity by selecting meaningful patch embeddings of varying sizes from the segmentation model. For instance, for the regional level caption generation, although the selected regions correspond to specific visual entities (e.g., the “airplane” in red boxes or the “person” in green boxes), the generated captions go beyond these isolated regions to describe broader scene elements, such as the “boarding ramp” or the “airport.”
@@ 3.1 Image Segmentation Module (ISM)
The first component of HBoP, ISM, selects patch embeddings ($E_{X}$) corresponding to image regions ($X=(X_{1},X_{2},...,X_{n})$) from the original image embeddings extracted using a Vision Transformer (ViT) Dosovitskiy et al. (2020) encoder. These regions are selected based on segmentation masks produced by a segmentation model. In our implementation, we use the Segment Anything Model (SAM) ***While we use SAM in our experiments, the HBoP framework is flexible and compatible with any segmentation model that can provide region masks. Additionally, several prior works Suo et al. (2023); Yu et al. (2023); Wang et al. (2025) have adopted training-free methods that incorporate unsupervised segmentation models for similar purposes. Additionally, a recent efficient implementation of SAM achieves up to 50× higher run-time speed, helping address concerns around computational overhead Zhao et al. (2023). Kirillov et al. (2023) due to its strong segmentation performance across diverse benchmarks. For a set of $p$ segmentation masks in the image, the resulting masks for the selected image regions would be: $M_{X}=\{\,M_{X_{1}},M_{X_{2}},...,M_{X_{p}}\,\}=\text{SAM}(X),X\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and channels of $X$.
@@ 3.2 Hierarchical Composition Module (HCM)
The global segmentation masks ($M_{G}$) are selected by choosing the top-$k$ (5 in our case) largest segmentation masks from $M_{X}$ after applying non-maximum suppression (NMS)NMS introduces a hyperparameter (IoU threshold), which is set to 0.1 in this case, to aggressively filter overlapping masks. While this step introduces a minor deviation from end-to-end processing, the hyperparameter is intuitive and fixed, requiring minimal tuning.Hosang et al. (2017):
$M_{G}$ $=\{M_{g_{1}},M_{g_{2}},...,M_{n_{g}}\},$
$M_{g_{i}}$ $=\text{NMS}(\text{Top-}k(M_{X})),\quad i=1,...,n_{g}$
@@
$M_{R}$ $=\{M_{r_{1}},M_{r_{2}},...,M_{K}\},$
$M_{r_{i}}$ $=\text{NMS}(\text{K-means}(M_{X})),i=1,...,K$
The hierarchical segmentation masks ($M_{g}$, $M_{r}$ and $M_{f}$) are used to extract relevant patch embeddings, $E_{g}$, $E_{r}$ and $E_{f}$ using $E_{X}$ from the first stage. We extract ($\odot$) the corresponding embeddings by concatenating the extracted patch embeddings of different levels (see Appendix C.1). Thus, the final selected image embeddings can be categorized as:
$E_{G}$ $=\{\,E_{g_{1}},E_{g_{2}},...,E_{g_{n_{g}}}\,\},E_{g_{i}}=E_{X}\odot M_{g_{i}}$
$E_{R}$ $=\{\,E_{r_{1}},E_{r_{2}},...,E_{K}\,\},E_{r_{i}}=E_{X}\odot M_{r_{i}}$
$E_{F}$ $=\{\,E_{f_{1}},E_{f_{2}},...,E_{n_{f}}\,\},E_{f_{i}}=E_{X}\odot M_{f_{i}}$
@@
Param SBERT $\uparrow$ CLIP-S $\uparrow$ PCD mBLEU-4 $\downarrow$ Div-2 $\uparrow$ SBERT $\uparrow$ CLIP-S $\uparrow$ PCD mBLEU-4 $\downarrow$ Div-2 $\uparrow$
@@
Table 1: Relevance and diversity scores across different models on the MSCOCO and Flickr30K datasets. HBoP achieves stronger diversity with higher Div-2 and PCD scores and a lower mBLEU-4 score compared to smaller VLMs and models trained to enhance diversity, while maintaining comparable relevance scores (SBERT and CLIP-S). Additionally, HBoP demonstrates competitive performance relative to much larger LLM-based VLMs. Cell colors indicate relative comparison to HBoP, with red showing higher values and blue showing lower values. Arrows next to each metric denote whether a higher (↑) or lower (↓) value indicates better performance.
3.3 Image Captioning Module (ICM)
$\text{HBoP}_{G}$ $=\{\,s_{g_{1}},...,s_{n_{g}}\,\},s_{g_{i}}=\text{BLIP}(E_{g_{i}})$
$\text{HBoP}_{R}$ $=\{\,s_{r_{1}},...,s_{K}\,\},s_{r_{i}}=\text{BLIP}(E_{r_{i}})$
$\text{HBoP}_{F}$ $=\{\,s_{f_{1}},...,s_{n_{f}}\,\},s_{f_{i}}=\text{BLIP}(E_{f_{i}})$
4 Results
HBoP achieves the best diversity scores while maintaining relevance among smaller VLMs. We evaluate the diversity and relevance of captions generated by different models in Table 1, using five captions per image. For HBoP, two global and three regional captions are sampledFine-grained captions are excluded from this evaluation because they function more as image tags than full descriptive captions.. Although HBoP increases the parameter count relative to BLIP, it remains significantly smaller than VLMs with LLMs, achieving a strong trade-off between diversity and model size. HBoP consistently achieves diversity scores closest to the gold-standard captions among smaller models, as measured by PCD (see Appendix C), mBLEU-4, and Div-2 Aneja et al. (2019b). Specifically, it reduces mBLEU-4 by over 60% and improves Div-2 by more than 30% compared to BLIP (NS) while maintaining comparable relevancy scores. We also compare our embedding-sampling approach to a baseline where segmented regions are directly cropped and captioned; while cropping improves diversity, it reduces relevance due to loss of global context. Full results are provided in Appendix C.
Compared to baselines such as BLIP Li et al. (2022a), Seq-CVAE Aneja et al. (2019b), and ModeCap§§§The dataset annotations and features necessary to train ModeCap are exclusively available for the MSCOCO dataset, making it difficult to replicate the experiments for fair comparison on the NoCaps and Flickr30k datasets.Chen et al. (2022), HBoP achieves the lowest mBLEU-4 and highest Div-2 scores. Notably, it even outperforms larger models like BLIP-2, Honeybee-7BCha et al. (2023), and LLaVA-1.5 Liu et al. (2023a) in several diversity metrics, despite using  4× to 13× fewer parameters. This highlights HBoP’s effectiveness as a lightweight alternative for generating diverse captions without the overhead of large-scale models.
HBoP maintains strong similarity between generated captions and reference texts and image-text alignment, as measured by SBERT and CLIP-Score, respectively. We use SBERT similarity as a practical proxy for human evaluation, as the reference captions are human-written; thus, high SBERT scores indicate strong semantic alignment and provide an approximate measure of caption quality from a human perspectiveThese high SBERT scores also imply that our framework is not negatively affected by the errors or failure cases of SAM. We also provide the performance and efficiency results of FastSAM Zhao et al. (2023) in Appendix D.4.. HBoP achieves scores comparable to BLIP, BLIP-NS, and LLaVA, while outperforming HoneyBee. Although BLIP-2 scores the highest, HBoP demonstrates a strong balance between relevance and diversity. Further semantic integrity evaluations are detailed in Appendix C.3.3.
5 Conclusion
6 Limitations
The current implementation of HBoP relies on bounding box approximations of segmentation masks to extract image embeddings. While effective, this may occasionally miss fine-grained or irregularly shaped image details. Exploring the use of full, irregular-shaped segmentation masks for embedding extraction is a promising direction for future work. Another limitation is that our approach primarily enhances factual diversity by focusing on distinct image regions, but it may be less effective for domains required for capturing cultural diversity Bayramli et al. (2025), as such interpretations often rely on external cultural knowledge beyond visual features.
7 Ethical Statement
Captions generated with HBoP might inadvertently contain harmful content. However, the final caption outputs mainly depend on the image content and the pretrained image captioning model. Therefore, unless the images themselves are harmful or the pretrained model produces unsafe captions, HBoP captions are expected to pose minimal risk.
8 Acknowledgement
This work was supported by Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)).
Appendix A Appendix
Appendix B Additional Related Works
B.1 Vision-Language Models (VLMs)
A growth of interest in VLMs has continued due to the wide availability of multimodal data on the web Waheed et al. (2025). Foundation VLMs can be applied to a range of tasks in a zero-shot manner. Notably, CLIP Radford et al. (2021b) jointly pre-trains an image encoder and a text encoder by maximizing and minimizing the cosine similarity of correct and incorrect image-text pair embeddings, respectively, with image-text contrastive (ITC) loss. In contrast, BLIP Li et al. (2022a) uses both ITC and image-text matching (ITM) loss for enhanced image-text data representation. Additionally, the BLIP Li et al. (2022a) captioner uses language modeling (LM) loss for autoregressive image caption generation along with a filter, capfilt, to improve the quality of image-text pairs for training.
Flamingo Alayrac et al. (2022) shows remarkable zero-shot ability in image captioning, visual question-answering (VQA), and image-text retrieval (ITR) tasks by leveraging the few-shot learning ability of pre-trained vision-only and language-only models. It simply interleaves input visual data with task-specific text examples, producing free-form texts for unseen visual data. Another general-purpose model, BEIT3 Wang et al. (2022b) with a Multiway Transformer structure, uses different types of modality experts to perform fusion and modality-specific training. A masked modeling objective on images only and image-text pairs is performed for computer vision tasks (e.g., image classification, semantic segmentation, object detection) and vision-language tasks (e.g., VQA), respectively. Whereas the VQA task uses a fused encoder for image-text pairs, the ITR task encodes images and texts independently with ITC loss. Lastly, sequence-to-sequence learning is applied to generate texts from images for the image captioning task. Inspired by these previous works, we propose a meta-VLM model that utilizes a pre-trained BLIP Li et al. (2022a) image captioning module to generate enhanced textual representations, which can later serve as useful data for various downstream tasks.
B.2 Hierarchical Representation
Identifying and extracting regions of interest within images is crucial for a hierarchical representation. The most intuitive way to achieve this would typically involve the use of object detectors Yao et al. (2019); Cornia et al. (2020); Zhang et al. (2021). However, the heavy computational demands of the object detectors inevitably lead to inefficiency during the inference stage Yao et al. (2019); Cornia et al. (2020); Zhang et al. (2021). In response, recent works sought to replace these cumbersome detectors by adopting visual concepts in the form of object tags Fang et al. (2022); Shukor et al. (2022) as an alternative. However, this detector-free approach is contingent upon the availability of object-specific data within the dataset. Employing pre-trained models is a more efficient way to identify areas of interest within images. GradCAM Selvaraju et al. (2017) highlights essential regions that the pre-trained models use to predict any target concept using its gradients with respect to feature map activations of the final convolutional layer. DINOv2 Oquab et al. (2023) capitalizes on existing self-supervised pre-trained models to generate robust, all-purpose visual features, supporting a wide array of tasks ranging from image-level classification to pixel-level segmentation. However, the image regions/features delineated by GradCAM/DINOv2 tend to show saliency for specific tasks and are unable to capture the full spectrum of visual representations. Conversely, SAM Kirillov et al. (2023) intricately segments every semantically significant component of an image into high-quality segmentation masks generated by prompting with various inputs such as point, box, mask, or free-form text, unrestricted by the types of tasks. In our framework, we integrate SAM Kirillov et al. (2023) to create semantically meaningful segmentation masks for an entire image automatically.
Several prior studies have incorporated the principles of hierarchy or multi-scale representation into their model architectures, aiming to enhance the alignment between images and texts Ji et al. (2021); Shao et al. (2023); Shukor et al. (2022). SHAN Ji et al. (2021) deconstructs the image-text matching process into two distinct facets: fragment-level and context-level alignments, enabling matches across three different scopes: local-to-local, global-to-local, and global-to-global. HiVLP Shao et al. (2023) leverages both low- and high-dimensional features to represent coarse and fine details. ViCHA Shukor et al. (2022) aligns images and texts across various layers of neural network encoders with the underlying assumption that each layer reflects varying semantic levels. Unlike these approaches, we divide the segmentation masks hierarchically and use the embeddings of the extracted individual image patches for caption generation.
B.3 Caption Evaluation
Common image captioning evaluation metrics, including BLEU Papineni et al. (2002a), METEOR Banerjee and Lavie (2005), ROUGE Lin (2004), and CIDEr Vedantam et al. (2015) scores, are primarily n-gram approaches that assess the quality of generated captions by considering their overlap with human-generated captions. Most SOTA VLMs frequently exhibit promising scores across these conventional evaluation metrics. However, these metrics are limited in their capabilities to measure the diversity of the generated captions. This limitation leads to a bias in these models towards generating an "average" and "safe" caption reflecting the most basic information in the image, rendering them less informative than human-generated captions. To address this gap, we incorporate several diversity metrics, including mBLEU-4, Div-2 Aneja et al. (2019b), and the proposed pairwise cosine distance (PCD), along with semantic integrity and relevance scores to ensure that the captions generated by our framework are not only diverse but also meaningful and directly relevant to the given image and human-annotated captions.
Appendix C Experiments
C.1 Implementation Details
The ISM (Section 3.1) employs the fully automated SAM with no prompting Kirillov et al. (2023), along with the image encoder initialized from ViT (ViT-L/16) pre-trained on ImageNet Dosovitskiy et al. (2021), following the same settings as BLIP Li et al. (2022a). Note that we use BLIP Li et al. (2022a) for captioning instead of BLIP-2 Li et al. (2023) since BLIP-2 uses intermediate representations trained on pairs of entire images and texts for caption generation using an LLM, which is not directly applicable to HBoP that uses pairs of image patches and texts.
The HCM (Section 3.2) creates the global level by selecting the top ($k=5$) masks with the largest areas and designating the remaining masks as fine-grained. To create the regional level, $K$-means clustering, with ($K=5$) clusters per image, is applied to the bounding boxes of the segmentation masks. For each cluster, we extract patch embeddings corresponding to the semantically segmented regions, projecting the clusters onto the image to identify relevant patches. The embeddings from each cluster are concatenated, and zero-padding is applied to preserve the original shape, ensuring that each cluster represents a semantically coherent region of the image. These regional embeddings are then used as input for caption generation in the captioning module. Lastly, the ICM (Section 3.3) follows the methodology outlined in Tiong et al., 2022.
While NMS with a threshold of 0.1 is applied at all three levels for all main experiments, we also conduct ablation experiments to examine the impact of varying the IoU threshold used in our sampling strategy. As shown in Table 2, while performance shows slight variation across different thresholds, our method consistently outperforms the baselines. This variance occurs because increasing the IoU threshold allows more overlapping image regions to be included, which tends to generate more repetitive captions and thus reduces diversity (PCD and Div-2 scores decreased).
IoU Threshold PCD Div-2 ↑
0.1 0.772 0.735
0.4 0.709 0.728
0.8 0.684 0.666
Table 2: Diversity results at different IoU thresholds.
All the model captions in Tables 1 and 6 are regenerated, except for Seq-CVAE Aneja et al. (2019b), where the results are taken directly from the original paper. While HBoP benefits from bounding box information, it is important to note that other baseline methods (e.g., ModeCap) have the additional advantage of explicit learning objectives to improve diversity. The exact prompts we use for Honeybee Cha et al. (2023) (top) and LLaVA-1.5/1.6 Liu et al. (2023a) are in Table 7.
C.2 Crop vs. Embedding Sampling Comparison
We present a comparison between our embedding-sampling approach (HBoP) and a baseline where segmented image regions are directly cropped and captioned using BLIP. Table 3 shows that while cropping can improve diversity scores, it often sacrifices relevance as indicated by SBERT. In contrast, HBoP preserves contextual understanding by sampling from full-image embeddings.
Models MSCOCO Flickr30k
SBERT ↑ mBLEU-4 ↓ Div-2 ↑ SBERT ↑ mBLEU-4 ↓ Div-2 ↑
BLIP (-NS) 56.00 1.00 0.179 55.78 1.00 0.179
BLIP (+NS) 57.23 0.66 0.387 46.99 0.66 0.387
Crop 52.03 0.10 0.600 50.00 0.08 0.610
HBoP 56.30 0.05 0.735 54.00 0.04 0.750
Table 3: Comparison of our embedding sampling approach (HBoP) with direct cropping of segmented regions.
C.3 Evaluation
C.3.1 Diversity
We measure the diversity in the generated captions using the cosine similarity between the sentence embeddings of all the corresponding captions per image. The comparison baselines are random captions, where each caption corresponds to different images, BLIP Li et al. (2022a) with and without nucleus sampling (NSUnless otherwise specified, all the BLIP models in this paper refer to BLIP with NS.) Holtzman et al. (2019), BLIP-2 Li et al. (2023), ModeCap Chen et al. (2022), Honeybee Cha et al. (2023), and gold captions******We exclude PnP-VQA since the captions are generated per question instead of per image, unlike other baselines.. The diversity of the generated captions ($s_{1},s_{2},...s_{n}$) per dataset instance††††††Note that $n=5$ for all dataset instances, and we use one global caption and five regional captions for HBoP. is measured using pairwise cosine distance (PCD):
In the above equation, coscos represents the cosine similarity of the input embeddings. We use sentence embeddings from a pre-trained sentence transformer model (all-MiniLM-L6-v2) Reimers and Gurevych (2019), denoted as $M$ in Eq. 1, that can capture the semantic relationships between captions An et al. (2024). This measure evaluates the extent to which the generated captions differ from each other per image. We report the final diversity score for each dataset as the average PCD score of all images in the dataset. Ideally, the PCD score should be lower than that of random captions that serve as the upper bound of the diversity score, but it should be higher than that for captions generated by existing baselines.
C.3.2 Relevancy
While confirming that each dataset contains captions with high semantic integrity is crucial, the captions must also be relevant to the corresponding images. We employ CLIP-Score Hessel et al. (2021) that calculates the correlation between visual and textual CLIP embeddings Radford et al. (2021b) using pre-trained ViT (openai/ clip-vit-base-patch32) without relying on human-generated references. Similar to the comparison baseline datasets for semantic integrity evaluation, we compare HBoP with PnP-VQA Tiong et al. (2022), BLIP Li et al. (2022a), BLIP-2 Li et al. (2023), gold captions, and random captions. We generate random captions by selecting five random captions for each image from a pool of HBoP captions corresponding to different images. In other words, although the random caption itself should make sense, they depict mismatched images. We randomly select one out of a total of five captions per image for each dataset and compute the correlation between CLIPScores of generated captions and gold captions.
C.3.3 Semantic Integrity
HBoP generates semantically meaningful captions. We evaluate the semantic integrity of HBoP captions using LLMs, LLaMA-2-13b Touvron et al. (2023) and GPT-4 Fu et al. (2023), which have shown high correlation with human judgment Chiang and yi Lee (2023); Liu et al. (2023b); Fu et al. (2023). Table 4 shows that HBoP achieves semantic integrity scores close to the gold captions and notably outperforms models like PnP-VQA Tiong et al. (2022). We attribute this improvement to our method’s ability to sample more meaningful image embeddings via the proposed Hierarchical Composition Module (HCM).
We prompt Llama-2-13B (Llama-2-13b-chat -hf) Touvron et al. (2023) to access the semantic integrity of HBoP captions along with gold and other baselines (PnP-VQA Tiong et al. (2022), BLIP Li et al. (2022a), BLIP-2 Li et al. (2023)) captions. Specifically, we randomly select two captions out of a total of five captions per image for each dataset and evaluate the semantic integrity by averaging the coherency and meaningfulness scores for each caption using the prompt shown in Table 8. We use the prompt "This is a picture of" to generate captions for all models in our experiments. This deliberate choice ensures a fair comparison of the general caption generation ability across models, as altering the prompt can yield significantly different results, making fair evaluation challenging.